OpenStudy (megannicole51):

Does the series converge or diverge? (n+ 1/n)^n

OpenStudy (megannicole51):

when I distribute the n does (1/n) become (1/n^n)?

OpenStudy (anonymous):

\[ \sum_n\left(\frac{n+1}{n}\right)^n \to \infty \]

OpenStudy (megannicole51):

awesome so it does diverge

OpenStudy (anonymous):

Yes. If you meant: \[ \sum_n\left(n+\frac{1}{n}\right)^n\to \infty \]It still diverges. When asking the question, could you put up what you got as an answer? Just so I don't get a warning for 'giving answers.' Thanks.

OpenStudy (megannicole51):

sure:)